# Event-Based Simulation of Stochastic Memristive Devices for Neuromorphic Computing

### v2

@@ Abstract
In this paper, we build a general modelling framework for memristors, suitable for the simulation of event-based systems such as hardware spiking neural networks, and more generally, neuromorphic computing systems. We propose an event-based modelling approach, extending and generalising an existing general model of memristors - the Generalised Metastable Switch Model (GMSM) [1]. In doing so we eliminate errors associated discrete time approximation, as well as offering potential improvements in terms of suitability for neuromorphic memristive system simulations. We then introduce a volatility state variable to allow for the unified understanding of disparate non-linear and volatile phenomena, including state relaxation, structural disruption, Joule heating, and non-linear drift in different memristive devices. We then demonstrate a practical implementation of the modelling framework, fit to a resistive drift dataset for titanium dioxide memristors, basing our fitting on a proposed linear conductance model for resistive drift in the devices. Finally, we highlight the application of the model to neuromorphic computing, through demonstrating the contribution of the volatility state variable to switching dynamics, resulting in frequency-dependent switching (for stable memristors acting as programmable synaptic weights) and the generation of action potentials (for unstable memristors, acting as spike-generators).
Index Terms:
memristors, neuromorphic computing, event-based models, volatility
@@ I Introduction
Figure 1: The event-based modelling framework. The framework consists of three components that are design choices: the Poisson rate parameter calculation for the state transitions, the connection between the exogenous inputs and the volatility state variable, and finally the readout equation for calculating the resistance from the state. The event-based modelling strategy is natively suited for event-driven systems and the volatility state variable can be used to explain and control key behaviours in a neuromorphic computing setting.
@@
The paper is structured as follows: in Section II, we begin by introducing our modelling framework, adapting an existing stochastic memristive model - the GMSM [1, 6]; we introduce an event-driven simulation approach, based on the evaluation of time-to-fire for multiple parallel Poisson point processes, while also allowing for continuous-time changes - such as changes in temperature or bias - to be incorporated. We then introduce the principle of the volatility state variable as a way to generalise and unify the description of a variety of nonlinear, temporary and autonomous evolution phenomena observed in different types of memristive device. Following the development of the model, in Section III, we demonstrate an example implementation of the modelling approach, applied to titanium dioxide memristors, using a generic Boltzmann occupation function to compute switching probabilities, modelling Joule heating using a volatility state variable, and proposing a linear readout equation motivated by a proposed physical model. We present a drift dataset collected for the titanium oxide memristors, and demonstrate the fit the model parameters to the observed data. Finally, in Section IV, we provide demonstrations of ways in which the proposed modelling framework can be used in neuromorphic simulations. We introduce the notion of stable and unstable memristors, and give illustrative examples of ways in which neuromorphic behaviours - namely, frequency-dependent switching for stable memristors and the emulation of action potentials for unstable memristors - can be brought about through an understanding of the effects of volatility on switching speed and the equilibrium point, and the subsequent appropriate selection of modelling parameters.
@@
II Event-Based Metastable Switching Model
In this section, we introduce the proposed modelling framework. In Section II-A we introduce an existing general and stochastic memristor model - the GMSM [1], detailing the introduction of a readout equation to separate state variables from summary parameters of interest such as the instantaneous device resistance. We then introduce our adaption to the event-based setting, basing the simulation on continuous time probabilistic point processes, rather than discrete time Bernoulli random variables in Section II. We introduce the “volatility state variable” in Section II-C to allow for unified description of a range of physical phenomena including autonomous state evolution, Joule heating, and nonlinear drift, which in turn will allow for the understanding of certain short term memory behaviours in memristive devices (see Section IV). In Section II-D, we specify the event-based simulation algorithm and a procedure for incorporating non-event-based inputs. We conclude the section with a discussion of the proposed model, in Section II-E.
II-A Modelling the Memristive State
II-A1 Generalised Metastable Switching Model
@@
Figure 2: The metastable switch model. The conductance is a linear function of the number of switches (parallel conducting channels) that are in the “on” state at any one time.
The GMSM bases its switching mechanism upon a Fermi-Dirac-like activation function (a logistic function), based on the temperature and the bias, which dictates the rate at which switching events occur. Two switching processes are modelled: low resistance to high resistance, and high resistance to low resistance. The probability of a single metastable switch transitioning from the low resistance state to the high resistance state in a given timestep $\Delta T$ is denoted by $P_{A}$, while the probability of a transition from the high resistance state to a low resistance state in $\Delta T$ is denoted by $P_{B}$ These two probabilities are given by
@@
where $\beta\triangleq\frac{q}{k_{B}T}$ is a physical parameter that introduces temperature-dependence and is the inverse of the thermal voltage ($k_{B}$ is Boltzmann’s constant and $q$ is the electronic charge), $\alpha=\frac{\Delta t}{t_{c}}$ is a ratio of the timestep to the characteristic timescale $t_{c}$ a parameter of the device.
@@
The GMSM models the device’s state evolution over time by considering the number of switches in either the high or low state and using the respective transition probabilities to define two binomial distributions that can be used to simulate the number of respective transitions of switches from low resistance to high, or vice versa for the second distribution, over a time-step. The normal approximation to a binomial random variable can be used, as $N$ (the number of metastable switches) becomes large, to develop a continuous stochastic model.
II-A2 State Variables
@@
In keeping with the terminology of the GMSM, we use the terms “switches”, or “metastable switches”, to refer to the state variables of the model (the volatility is separately dealt with). For the purposes of this work, we assume that both the number of state variables ($N$) and the multiplicity of the individual state spaces are finite. Specifically, we consider the case where $k=2$ i.e. binary switches. Thus, the multiplicity of the state space (number of possible memristive states) is $2^{N}$. We use $n$ to denote the number of switches in the state that is conducive to conduction, which we call the “up state”, the “high conductance state”, or the “low resistance state”. Then, $m=N-n$ denotes the number of switches in the state that is not conducive to conduction, which we call the “down state”, the “low conductance state”, or the “high resistance state”. Switching from a lower to a higher resistance state is termed “up” switching, while switching in the opposite direction is termed “down” switching, reflecting the respective increase or decrease in the resistance. Note that due to the deterministic relationship between $m$ and $n$, there is effectively only one state variable, yet we make use of both for notational convenience. It has been shown that a single state variable is sufficient to explain the behaviour of various memristive devices [9, 10].
II-A3 Readout Equation
Note that the voltage-current relationship for a memristive device may not always be instantaneously linear. Indeed, an exponential or $\sinh$ relationship is possible depending on the nature of the device being considered [1]. Different behaviour can be engineered by relating the memristive state variable, $n$, to a separate quantity such as the parameter for a Shockley diode equation, or the scaling factor for a $\sinh$ voltage-current relation.
Rather than requiring that the resistance be linearly linked to the state variable, we instead use a readout equation to provide flexibility in simulation. The readout equation takes as input the memristive state variable $n$, and outputs a resistance. This allows for a consistent description of the state, and allows for a modular approach that separates state evolution dynamics from the resistance, which can be considered a summary statistic of a more complex state. In Section III, we see an example of a thresholding, non-linear readout equation, that links the state variable to the voltage-current characteristics in a non-linear way. The role of the readout equation in the framework is visualised in Figure 1.
II-B Event-Based Switching and the Rate Parameter Model
@@
II-C Volatility State Variable
@@
    Accelerated switching behaviour, where the frequency or duration (spacing in time) of applied voltage pulses impacts the rate of switching (see Section IV-C1).
@@
II-D Practical Simulations
@@
II-D1 Point Process Sampling
Considering the cumulative distribution function for this exponential distribution, we can use use inverse transform sampling (the Smirnov transformation) to generate samples for the point process:
@@ II-D1 Point Process Sampling
II-D2 Changes to the Distribution at Events
@@
We use the memoryless property of the exponential distribution, which states that if an event has not occurred by time $\tau$ then the distribution for an event to occur at time $t+\tau$ is still exponentially distributed. The only distinction is that we now modify the rate parameter for each distribution, which can be calculated as a simple linear scaling of the rate parameter, since the sum of N independent and identically distributed exponential distributions with rate parameter $\lambda$ is an exponential distribution with rate parameter $N\lambda$, as described in Section II-B.
II-D3 Simulation Algorithm
@@
Note that in the case the value of $n$ falls to zero, or reaches $N$, we simply wait for an appropriate event to occur, to increase or decrease the value of $n$, respectively.
@@
In order to incorporate non-event based inputs, including variables that may be continuous functions of time such as temperature or applied bias, or piecewise constant functions, we can sample these using threshold-based, or event-driven, sampling. For example, the volatility state variable, introduced in Section II-C, is a continuous function of time. Given an independent input signal $m_{i}(t)$, we can define the associated marked point process, defined by the set of times $\mathcal{T}_{i}$, at which the events occur, to be:
$\mathcal{T}_{i}$ :=$:=\{t:f_{i}(t)\in\mathcal{M}_{i}\}$
@@
where, for continuous functions the set $\mathcal{M}_{i}$ contains the function values, changing in steps of $a$:
$\mathcal{M}_{i}$ =$=\{f(0),f(0)+a,\cdots,\lim_{k\rightarrow\infty}f(0)+ka\}$
$\cup\{f(0),f(0)-a,\cdots,\lim_{k\rightarrow\infty}f(0)-ka\}$
In contrast, for piecewise constant (discontinuous) functions, $\mathcal{M}_{i}$ will simply be the set of associated amplitudes at times $\mathcal{T}_{i}$, known ahead of time.
Due to the memoryless nature of the Poisson process, incorporating such sampling into the model is straightforward - whenever $t\in\mathcal{T_{i}}$, for a particular associated set of events $i$, we enforce the event and then resample times to firing for the Poisson switching processes. The canonical example of an input that can be modelled in this way, is the driving input voltage, that takes on predefined values at predetermined points in time.
@@
Although the method of conversion for continuous functions for inclusion in the event-based modelling framework works to minimise the error associated with changes to parameters affecting the rate equations not matching those used during the simulation - which are dependent purely on external stimuli - additional challenges are presented in the case of variables that are instead, or also, functions of the device state themselves. For example, if Joule heating is dependent on the current through the memristor, then this variable must be re-evaluated whenever there is a change in the resistance, or in the voltage. As such, for variables in the rate equations that are functions of both external stimuli and the internal state variable, these should be re-evaluated every time any event happens, so as to avoid errors associated with the drift of the independent variables of the function away from the values used for the calculation.
We can incorporate these variables by considering them to be a subset of the variables by allowing the updates to be governed by function evaluation at pre-defined times, $\mathcal{T}_{i}$, as described previously, in order to account for external (autonomous) changes to the variable, in addition to at any point when a dependent variable of the function changes. Let $\{j_{0}\cdots j_{j}\}\in\{m_{0}\cdots m_{i}\}$ be the set of functions which depend on at least one other variable. We note that the up and down switching rates themselves, $\lambda_{\text{up}}(n)$ and $\lambda_{\text{down}}(n)$, are within the subset $\{j_{0}\cdots j_{j}\}$, since they are updated whenever the variable $n$ is modified.
The algorithm for switching, including the incorporation of additional updates to parameters which are dependent on external stimuli - not limited to state switching events - is given in Algorithm 1.
Algorithm 1 Event-based modelling procedure, incorporating additional non-event-based inputs.
1:$R\leftarrow R_{\text{init}}$
2:$n\leftarrow n(R)$
3:$M_{i}\leftarrow\mathcal{M}_{i}[0]$
4:$\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(\{M_{i}\})$
5:$\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(\{M_{i}\})$
6:$c_{i}\leftarrow 0$
7:$t\leftarrow 0$
8:$F\leftarrow 0$triangleright State variable which denotes whether a firing event has occurred.
9:while $t\leq\Delta T$ do
10:     $t_{\text{down}}\sim\text{Exp}((N-n)\cdot\lambda^{\prime}_{\text{down}})$
11:     $t_{\text{up}}\sim\text{Exp}(n\cdot\lambda^{\prime}_{\text{up}})$
12:     for all $i$ do $\triangleright$ Firing for the state-independent variables
13:         if $F=0\;\textbf{and}\;c_{i}\leq|\mathcal{T}_{i}|\;\textbf{and}\;\mathcal{T}_{i}[c_{i}+1]\leq\{t_{\text{up}},t_{\text{down}},\mathcal{T}_{k}[c_{k}+1]\;\forall k\neq i\}$ then
14:              $c_{i}\leftarrow c_{i}+1$
15:              $t\leftarrow\mathcal{T}_{i}[c_{i}]$
16:              $M_{i}\leftarrow m_{i}({M}_{i}[c_{i}],n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
17:              $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
18:              $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
19:              $F\leftarrow 1$
20:     if $F=0\;\textbf{and}\;(t_{\text{down}}\leq t_{\text{up}}\;\textbf{or}\;n=N)\;\textbf{and}\;n\neq 0$ then
21:         $n\leftarrow n-1$
22:         $t\leftarrow t+t_{\text{down}}$
23:         for all $j$ do $\triangleright$ Update the state-dependent variables
24:              $J_{i}\leftarrow j_{i}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
25:              $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
26:              $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
27:     else if $F=0\;\textbf{and}\;(t_{\text{up}}<t_{\text{down}}\;\textbf{or}\;n=0)$ then
28:         $n\leftarrow n+1$
29:         $t\leftarrow t+t_{\text{up}}$
30:         for all $j$ do
31:              $J_{i}\leftarrow j_{i}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
32:              $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
33:              $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
34:     $R\leftarrow R(n)$
35:     $F\leftarrow 0$
@@
If necessary, we can convert non-uniform series from the event-based model to uniform data in time through sampling, albeit introducing errors associated with the discretisation in time, as shown in Algorithm 2. $\mathcal{T}_{switch}$ denotes the ordered set of switching times, and $\mathcal{R}$ denotes the associated ordered set of values of a state variable at the times $\mathcal{T}_{switch}$. $T_{\text{sample}}$ denotes the sampling period and $T_{\text{tot}}$ denotes the maximum desired time to sample for.
@@
$i\leftarrow-1$
2:$t_{\text{curr}}\leftarrow 0$
3:$\mathcal{R}_{\text{sample}}\leftarrow\{\}$
@@
16:$\textbf{return}\;\mathcal{R}_{\text{sample}}$
II-E Discussion
II-E1 Impact of the Parameter $N$
One strength of a discrete modelling strategy, with a finite number of possible system states (in our case, $N$), is that simulations can be conducted at different resolutions, depending on precision requirements, and the scale of interest. For example, if a memristive device is to be used as a binary storage device, where only the states $1$ and 00 are distinct, then a value of $N=2$ suffices to model the readout value of interest. In contrast, for a simulation where multiple stable states are to be for information storage, a larger value of $N$ may be required. The “best” value of $N$ to use is a matter of perspective - it depends on the extent to which regions of device resistance or behaviour are considered to be equivalent for the intended purpose, as exemplified by binary storage, where the many possible configurations of the device are summarised by a single bit (2 possible values) of information.
As a consequence of the fact that the variance and mean of a Poisson process are equal to the rate of the process, the magnitude of the parameter $N$ will scale the expected value and the variance of the final state in a linear fashion. As a result, the standard deviation in the final state, after a given delay, is proportional to $\sqrt{N}$. We can therefore modulate stochasticity by by changing $N$. When we increase N, the average jump time decreases linearly with N (the rates of up and down jumps increase proportional to N). This means that over the timescale of interest, many more events will tend to occur, with each trial also contributing to a smaller percentage change in the resistance, resulting in a decrease in the variance with increasing $N$. Continuous, deterministic switching behaviour can be approximated in the limit as $N\rightarrow\infty$.
II-E2 Elimination of Discrete-Time Sampling Error
We note that the approximation in Equation 1 is valid only for small values of $\Delta t$. If we let the value of $\Delta t$ grow large, then errors associated with discretising the timestep begin to emerge. The larger the value of $\Delta t$, the more significant a change will have occurred in the parameters used to calculate the transition probabilities, leading to modelling inaccuracies. This is a problem for any discrete-time approximation of a continuous-time problem i.e. any finite difference method for approximating a continuous-time differential equation. Error-free modelling can only be achieved in the limit as the timestep vanishes, i.e., $\Delta t\rightarrow 0$. The switching probability at each point in time will change with each switching event; however, updates are not instantaneously incorporated into the model, but are only applied every $\Delta t$ seconds, over which time many switching events may have occurred, resulting in an error. Such errors compound over time.
An advantage of our event-based adaptation of the GMSM is to eliminate the error associated with discrete time-stepped simulation. However, it should be noted that using a value of N that is too coarse will still introduce a quantisation error. This error does not, however, compound over time and the expected value of the process after given number of switching events will not be biased, since the value of $n_{eq}$ scales linearly with $N$, and the associated parameters of the readout equation can be adjusted accordingly.
Note that integrate and fire sampling for continuous functions of time, as described in Section II-D3, will unavoidably introduce discretisation errors. If the functions vary slowly in time, then the number of additional events due to the introduction of additional functions will be small. However, if the functions vary quickly in time and the number of generated events is large, then the benefits of event-based modelling approaches diminish, and the model’s events become dense in time. This still, however, offers benefits over a uniform sampling approach, since it associates the density of samples in time with the rate of information change that impact the model updates (and hence the error associated with discrete-time sampling).
III Metal Oxide Memristor Modelling

### v3

@@ Abstract
In this paper, we build a general modelling framework for memristors, suitable for the simulation of event-based systems such as hardware spiking neural networks, and more generally, neuromorphic computing systems composed of three independent components: i) an event-based modelling approach, extending and generalising an existing general model of memristors - the Generalised Metastable Switch Model (GMSM) [1] - eliminating errors associated with discrete time approximation, as well as offering potential improvements in terms of suitability for neuromorphic memristive system simulations; ii) a volatility state variable to allow for the unified understanding of disparate non-linear and volatile phenomena, including state relaxation, structural disruption, Joule heating, and non-linear drift in different memristive devices; and iii) a readout equation that separates the latent state variable evolution from explicit variables of interest such as an instantaneous resistance. We exhibit an illustrative implementation of this framework, fit to a resistive drift dataset for titanium dioxide memristors, based on a proposed linear conductance model for resistive drift in the devices. Finally, we highlight the application of the model to neuromorphic computing, through demonstrating the contribution of the volatility state variable to switching dynamics, resulting in frequency-dependent switching (for stable memristors acting as programmable synaptic weights) and the generation of action potentials (for unstable memristors, acting as spike-generators).
@@ I Introduction
Figure 1: The event-based modelling framework. The framework consists of three components that are design choices: the Poisson rate equation (for state transitions), the volatility equation(s) (linking the exogenous inputs to the volatility state variable(s)), and the readout equation (that maps an internal state to a visible quantity such as the resistance).
@@
The paper is structured as follows: in Section II, we begin by introducing an existing stochastic memristive model - the GMSM [1, 6] and the state modelling approach. We introduce our modelling framework in Section III, composed of three main components: event-driven simulation, based on the evaluation of time-to-fire for multiple parallel Poisson point processes; volatility state variables as a way to generalise and unify the description of a variety of nonlinear, temporary and autonomous evolution phenomena observed in memristors; and the notion of the readout equation to allow for the separation of state variables from visible device readings such as resistance. Following the introduction of the modelling framework, in Section IV, we demonstrate an example implementation, applied to titanium dioxide memristors. We present a drift dataset collected for the titanium oxide memristors, and demonstrate the fit of the model parameters to the observed data. In Section V, we offer a discussion of event-based component of the modelling approach, evaluating it qualitatively and comparing it to discrete timestep approximations. Finally, in Section VI, we provide demonstrations of ways in which the proposed modelling framework can be used in neuromorphic simulations. We introduce the notion of stable and unstable memristors, and give illustrative examples of ways in which neuromorphic behaviours - namely, frequency-dependent switching for stable memristors and the emulation of action potentials for unstable memristors - can be brought about through an understanding of the effects of volatility on switching speed and the equilibrium point, and the subsequent appropriate selection of modelling parameters. The source code for our model and simulations is made publicly available online [7].
@@
II Background
@@
Figure 2: The metastable switch model. The conductance is a linear function of the number of switches (parallel conducting channels) that are in the “on” state (blue - S2 and S4 above) at any one time, as opposed to those in the “off” state (red - S1 and S3 above).
The GMSM bases its switching mechanism upon a Fermi-Dirac-like activation function (a logistic function), based on the temperature and the bias, which dictates the rate at which switching events occur. Two switching processes are modelled: low resistance to high resistance, and high resistance to low resistance. The probability of a single metastable switch transitioning from the low resistance state to the high resistance state in a given timestep $h$ is denoted by $P_{A}$, while the probability of a transition from the high resistance state to a low resistance state is denoted by $P_{B}$ . These two probabilities are given by
@@
where $V$ is the applied bias, $V_{A}$ and $V_{B}$ can be thought of as energy barrier heights that modulate the switching probabilities, $\beta\triangleq\frac{q}{k_{B}T}$ is a physical parameter that introduces temperature-dependence and is the inverse of the thermal voltage ($k_{B}$ is Boltzmann’s constant and $q$ is the electronic charge), and $\alpha=\frac{h}{t_{c}}$ is a ratio of the timestep to the characteristic timescale $t_{c}$ - a parameter of the device.
@@
The GMSM models the device’s state evolution over time by considering the number of switches in either the high or low state and using the respective transition probabilities to define two binomial distributions that can be used to simulate the number of respective transitions of switches from low resistance to high, or vice versa for the second distribution, over a time-step. The normal approximation of binomial random variables can be used, as $N$ becomes large, to form a continuous stochastic model.
@@
In keeping with the terminology of the GMSM, we use the terms “switches”, or “metastable switches”, to refer to the state variables of the model (the volatility is separately dealt with). For the purposes of this work, we assume that both the number of state variables ($N$) and the multiplicity of the individual state spaces are finite. Specifically, we consider the case where $k=2$ - i.e. binary switches. Thus, the multiplicity of the state space (number of possible memristive states) is $2^{N}$. We use $n$ to denote the number of switches in the state that is conducive to conduction, which we call the “up state”, the “high conductance state”, or the “low resistance state”. Then, NnN-n denotes the number of switches in the state that is not conducive to conduction, which we call the “down state”, the “low conductance state”, or the “high resistance state”. Switching from a lower to a higher resistance state is termed “up” switching, while switching in the opposite direction is termed “down” switching, reflecting the respective increase or decrease in the resistance. It has been shown that a single state variable is sufficient to reasonably explain the behaviour of various memristive devices [10, 11].
III Modelling Framework
In this section, we introduce the proposed modelling framework. Our framework consists of three modular components: a rate equation, volatility equation(s), and a readout equation, which can each be different depending on the memristive device being modelled. In Section III-A, we start by introducing our adaption to the event-based setting, basing the simulation on continuous time probabilistic point processes, rather than discrete time Bernoulli random variables. Next, in Section III-B, we introduce the “volatility state variable” to allow for unified description of a range of physical phenomena including autonomous state evolution, Joule heating, and nonlinear drift, which in turn will allow for the understanding of certain short term memory behaviours in memristive devices (see Section VI). Section III-C introduces the readout equation to separate state variables from summary parameters of interest such as the instantaneous device resistance. In Section III-D, we specify the event-based simulation algorithm and a procedure for incorporating non-event-based inputs.
III-A Event-Based Switching and the Rate Parameter Model
@@
III-B Volatility State Variable
@@
    Accelerated switching behaviour, where the frequency or duration (spacing in time) of applied voltage pulses impacts the rate of switching (see Section VI-C1).
@@
III-C Readout Equation
The voltage-current relationship for a memristive device need not always be instantaneously linear. Indeed, an exponential or $\sinh$ relationship is possible depending on the nature of the device being considered [1]. Different behaviour can be engineered by relating the memristive state variable, $n$, to a separate quantity such as the parameter for a Shockley diode equation, or the scaling factor for a $\sinh$ voltage-current relation, depending on the device.
Rather than requiring that the resistance be linearly linked to the state variable, we can instead use a readout equation to provide flexibility in simulation. The readout equation takes as input the memristive state variable $n$, and outputs a resistance. This allows for a consistent description of the state, and allows for a modular approach that separates state evolution dynamics from the resistance, which can be considered a summary statistic of a more complex state. In Section IV, we see an example of a thresholding, non-linear readout equation, that links the state variable to the voltage-current characteristics in a non-linear way. The role of the readout equation in the framework is visualised in Figure 1.
III-D Practical Simulations
@@
III-D1 Point Process Sampling
Considering the cumulative distribution function for this exponential distribution, we can use inverse transform sampling (the Smirnov transformation) to generate samples for the point process:
@@ III-D1 Point Process Sampling
III-D2 Changes to the Distribution at Events
@@
We use the memoryless property of the exponential distribution, which states that if an event has not occurred by time $\tau$ then the distribution for an event to occur at time $t+\tau$ is still exponentially distributed. The only distinction is that we now modify the rate parameter for each distribution, which can be calculated as a simple linear scaling of the rate parameter, since the sum of N independent and identically distributed exponential distributions with rate parameter $\lambda$ is an exponential distribution with rate parameter $N\lambda$, as described in Section III-A.
III-D3 Simulation Algorithm
@@
Note that in the case the value of $n$ falls to zero, or reaches $N$, we simply wait for an appropriate event to occur, to increase or decrease the value of $n$, respectively. Algorithm 1 shows the switching process for a total simulation time $T_{\text{max}}$.
@@
In order to incorporate non-event based inputs, including variables that may be continuous functions of time such as temperature or applied bias, or piecewise constant functions, we can choose to apply one of two approaches. If the values are unknown beforehand, we can include them as additional point processes with deterministic interarrival times equal to a fixed sampling timestep, $t_{\text{sample}}$, which lower bounds the error in their values to that of a discrete time simulation, introducing a point process, $\mathcal{T}_{\text{sample}}$, for updates, where e.g. a Euler integration update step is performed:
$\mathcal{T}_{\text{sample}}=\{t_{\text{sample}},2t_{\text{sample}},3t_{\text{sample}}\ldots,\}$
@@
However, if signal values are known beforehand then we can also sample using threshold-based, or event-driven, sampling. Given an independent input signal $m(t)$, we can define the associated marked point process, with the set of times $\mathcal{T}$, at which the events occur, to be:
$\mathcal{T}$ $=\{t:m(t)\in\mathcal{M}\}$
@@
where, for continuous functions the set $\mathcal{M}$ contains the function values, changing in steps of $\Delta m$:
$\mathcal{M}_{i}$ $=\{m(0),m(0)+\Delta m,\cdots,m(0)+l\Delta m\}$
$\cup\{m(0),m(0)-\Delta m,\cdots,m(0)-l\Delta m\}$
i.e. a set of signal values at specified boundaries and their associated event times. For piecewise constant (discontinuous) functions, $\mathcal{M}$ will simply be the set of amplitudes at points when the function changes, $\mathcal{T}$. Due to the memoryless nature of the Poisson process, incorporating such sampling into the event-based modelling is straightforward - whenever $t\in\mathcal{T}$, for a particular associated set of events, we run the event (update all continuous time variables and rate equations) and resample times to fire for the Poisson switching processes.
@@
Although the method of conversion for continuous functions for inclusion in the event-based modelling framework works to minimise the error associated with changes to parameters affecting the rate equations not matching those used during the simulation - which are dependent purely on external stimuli - additional challenges are presented in the case of variables that are instead, or also, functions of the device state themselves. For example, volatility functions that update volatility state variables, such as Joule heating, will likely be dependent on the current through or voltage across the memristor. In this case, the variable must be re-evaluated whenever there is a change in the resistance, or in the voltage. Indeed, for rate equations that are functions of both external stimuli and the internal state variable, these should be re-evaluated every time any event happens, so as to avoid errors associated with the drift of the independent variables of the function away from their original values. Let $\{j_{0}\cdots j_{l}\}$ be the update functions for the variables $\{J_{0}\cdots J_{l}\}$, each with initial values $J_{l}^{0}$, which depend on at least one other variable for their updates and as such cannot be marked processes since they cannot be precomputed. A special marked process $\mathcal{T}_{\text{sample}}$ is associated with this. These are updated whenever any other variable $n$ is modified (i.e. an event occurs) by e.g. Euler integration.
Algorithm 1 Event-based modelling procedure.
[…]
5:while $t\leq T_{\text{max}}$ do
6:  $t_{\text{down}}\sim\text{Exp}((N-n)\cdot\lambda^{\prime}_{\text{down}})$
7:  $t_{\text{up}}\sim\text{Exp}(n\cdot\lambda^{\prime}_{\text{up}})$
8:  for all $i$ do $\triangleright$ State-independent firing.
9:   if $F=0\;\textbf{and}\;c_{i}\leq|\mathcal{T}_{i}|\;\textbf{and}\;\mathcal{T}_{i}[c_{i}]-t\leq\{t_{\text{sample}},t_{\text{up}},t_{\text{down}},\mathcal{T}_{k}[c_{k}]-t\;\forall k\neq i\}$ then
10:     $t_{\text{diff}}\leftarrow\mathcal{T}_{i}[c_{i}]-t$
11:     for all $j$ do
12:      $J_{l}\leftarrow j_{l}(t_{\text{diff}},n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
13:     $M_{i}\leftarrow\mathcal{M}_{i}[c_{i}]$
14:     $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
15:     $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
16:     $t\leftarrow\mathcal{T}_{i}[c_{i}]$
17:     $c_{i}\leftarrow c_{i}+1$
18:     $F\leftarrow 1$
19:  if $F=0\;\textbf{and}\;(t_{\text{down}}<t_{\text{sample}})\;\textbf{and}\;(t_{\text{down}}\leq t_{\text{up}}\;\textbf{or}\;n=N)\;\textbf{and}\;n\neq 0$ then
20:   for all $j$ do
21:     $J_{l}\leftarrow j_{l}(t_{\text{down}},n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
22:   $n\leftarrow n-1$
23:   $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
24:   $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
25:   $t\leftarrow t+t_{\text{down}}$
26:  else if $F=0\;\textbf{and}\;(t_{\text{up}}<t_{\text{sample}})\;\textbf{and}\;(t_{\text{up}}<t_{\text{down}}\;\textbf{or}\;n=0)$ then
27:   for all $j$ do
28:     $J_{l}\leftarrow j_{l}(t_{\text{up}},n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
29:   $n\leftarrow n+1$
30:   $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
31:   $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
32:   $t\leftarrow t+t_{\text{up}}$
33:  else if $F=0$ then
34:   for all $j$ do
35:     $J_{l}\leftarrow j_{l}(t_{\text{sample}},n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
36:   $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
37:   $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{l})$
38:   $t\leftarrow t+t_{\text{sample}}$
39:  $R\leftarrow f_{ro}(n)$
40:  $F\leftarrow 0$
@@
If necessary, we can convert non-uniform series from the event-based model to uniform data in time through sampling, albeit introducing errors associated with the discretisation in time, as shown in Algorithm 2. $\mathcal{T}_{switch}$ denotes the ordered set of switching times, and $\mathcal{R}$ denotes the associated ordered set of values of a state variable at the times $\mathcal{T}_{switch}$. $T_{\text{sample}}$ denotes the sampling period and $T_{\text{tot}}$ denotes the maximum desired time to sample.
@@
$1i\leftarrow$-1
[…]
@@
[…]
IV Metal Oxide Memristor Modelling